umerlodhi (umerlodhi):

A problem states: "There are 7 more girls than boys in a class. There are 23 students in the class in all. How many girls are there in the class?" What are the unknowns in this problem? A. the number of boys and the number of girls in the class B. the number of students in the class C. the number of boys in the class D. the number of girls in the class

OpenStudy (anonymous):

c

umerlodhi (umerlodhi):

ok i have one more

umerlodhi (umerlodhi):

Zach is 4 years older than twice his sister Maya’s age. Let m represent Maya’s age. Which expression represents Zach’s age? A. 2m + 4 B. m + 4 C. 4m + 2 D. 2m

umerlodhi (umerlodhi):

Solve the given problem. Melanie is 2 years older than twice her brother Bill’s age. If Bill is 5 years old, how old is Melanie? A. 4 B. 10 C. 12 D. 14

OpenStudy (anonymous):

its c my bad

OpenStudy (anonymous):

d.14

umerlodhi (umerlodhi):

what

umerlodhi (umerlodhi):

the second one is c

umerlodhi (umerlodhi):

The area of a rectangle is 70 square inches. The length of the rectangle is 3 inches longer than the width. Which equation models this situation? A. 3w = 70 B. w + 3 = 70 C. w + 3w = 70 D. w(w + 3) = 70

umerlodhi (umerlodhi):

The area of a rectangle is 70 square inches. The length of the rectangle is 3 inches longer than the width. Which equation models this situation? A. 3w = 70 B. w + 3 = 70 C. w + 3w = 70 D. w(w + 3) = 70

OpenStudy (anonymous):

yea the second one c... d w(w+3)=70